OpenStudy (anonymous):

how do I graph the following equations y=3x+6, y=x-2, y=3x+10

OpenStudy (anonymous):

You can find the intercepts and make your life easier

OpenStudy (anonymous):

At that point it would be just connecting the dots

OpenStudy (anonymous):

To find x intercept; let y = 0 and solve for x. To find y intercept; let x = 0 and solve for y.

OpenStudy (anonymous):

How? Its pretty difficult

OpenStudy (anonymous):

so the first one is y = 3x+6, find the x intercept, 0=3x+6, x = -6/3 = -2 y intercept is y = 6